One kg of air is compressed from 1m cube to 0.5m cube. What will be the effect on density of air?
Answer #1

density = mass / volume

if mass = 1 kg = constant =

density is inversely proportional to volume


d1 / d2 = V2 / V1

d1 / d2 = 0.5 / 1


d2 = d1 / 0.5

d2 = 2 d1............... doubles